Good evening - Can someone help me with the following equation?
Find the x-intercepts
y = x^2-5x-10
Thank you so much!

Answer by stanbon(75887)

Find the x-intercepts
y = x^2-5x-10
To find the x-intercepts, let y=0, then solve for "x".
x^2-5x-10=0
Using the quadratic formula get:
x=[5+sqrt(25-4(1)(-10)]/2 = [5+sqrt65]/2
OR
x=[5-sqrt65]/2
Cheers,
Stan H.